OpenStudy (anonymous):

Solve 81x = 27 ^x + 2.

OpenStudy (anonymous):

Note that 81 = 3^4 and 27 = 3³. Then, we have.. 3^(4x) = 3^(3(x + 2)) 3^(4x) = 3^(3x + 6) Set the exponents equal to each other.. 4x = 3x + 6 Subtract both sides by 3x... 4x - 3x = 3x - 3 + 6 So x = 6 [D] I hope this helps!